If the mean and the standard deviation of a continuous random variable that is normally distributed are 20 and 5, respectively, find an interval that contains 68% of the distribution.

a. 18,24

b. 12,25

c. 10,30

d. 15,25
Answer #1

Given Mean u =20 and standard deviation o=5 Given 68% of the distribution a=1-0.68=0.32 a/2=0.16 z intervals for the middle 6
Answer #2
If the mean and the standard deviation of a continuous random variable that is normally distributed are 28 and 3 respectively, using the Empirical rule find an interval that contains 95% of the distribution
answered by: Amanda